OpenStudy (anonymous):

CAN I PLEASE JUST GET HELP!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!! Create a unique example of dividing a polynomial by a monomial and provide the simplified form. Explain, in complete sentences, the two ways used to simplify this expression and how you would check your quotient for accuracy.

OpenStudy (anonymous):

We have to divide the two x2+3x+2x+1 Let's first simplify this by long division x _________________________________ x+1 | x2+3x+2 -(x2+x) x+2 Next step x+2 _________________________________ x+1 | x2+3x+2 -(x2+x) 2x+2 -( 2x+2 ) ___________________________ 0 So we get x2+3x+2x+1=x+2

OpenStudy (anonymous):

I think your answer is x+7

OpenStudy (anonymous):

Hope it helped!

OpenStudy (anonymous):

@ganeshie8 can you help me with this please

ganeshie8 (ganeshie8):

krisylynn example is looking good :)

OpenStudy (anonymous):

so it is correct?

ganeshie8 (ganeshie8):

It is correct

OpenStudy (anonymous):

thankx man

ganeshie8 (ganeshie8):

np, read it like below :- We have to divide the two x2+3x+2 / x+1 Let's first simplify this by long division x _________________________________ x+1 | x2+3x+2 -(x2+x) x+2 Next step x+2 _________________________________ x+1 | x2+3x+2 -(x2+x) 2x+2 -( 2x+2 ) ___________________________ 0 So we get x2+3x+2 / x+1=x+2